# Limitations of Data-Driven Spectral Reconstruction – Optics-Aware Analysis and Mitigation

### v2

@@ 3.3. Modeling metamerism
Interestingly, metamer generation from existing spectra has been studied in color science and spectral rendering to accurately model the scenes using various methods, e.g., metameric black (van Trigt, 1994; Finlayson and Morovic, 2005; Alsam and Lenz, 2007) and spectral uplifting (Jakob and Hanika, 2019; Van De Ruit and Eisemann, 2023; Belcour et al., 2023). In this work, to support our analysis, we adopt the metameric black approach to generate metamers, whereas we note that other metamer generation methods can also be employed for the same purpose.
A spectrum ${\mathbf{S}}$ can be projected onto two orthogonal subspaces, one for the fundamental metamer ${\mathbf{S}}^{*}$, and the other for metameric black ${\mathbf{B}}$ (Cohen and Kappauf, 1982; Viénot and Brettel, 2014). A new metameric spectrum ${\mathbf{S}}^{\prime}$ is then
(3) ${\mathbf{S}}^{\prime}={\mathbf{S}}^{*}+\alpha{\mathbf{B}},$
@@
3.4. Performance evaluation metrics
Consider a hyperspectral image ${\mathbf{X}}_{k,i,j}$ and its estimate $\hat{{\mathbf{X}}}_{k,i,j}$, where $k$ is the spectral index, and $i,j$ are spatial indices. The reconstruction quality can be evaluated in various ways. The NTIRE 2022 spectral reconstruction challenge (Arad et al., 2022) adopts two numerical metrics, Mean Relative Absolute Error (MRAE),
(4) ${\mathrm{\mathbf{MRAE}}}=\frac{1}{KMN}\sum_{k,i,j}\left|\hat{{\mathbf{X}}}_{k,i,j}-{\mathbf{X}}_{k,i,j}\right|/{\mathbf{X}}_{k,i,j},$
@@ 3.4. Performance evaluation metrics
(6) ${\mathrm{\mathbf{RMSE}}}=\sqrt{\frac{1}{KMN}\sum_{k,i,j}\left(\hat{{\mathbf{X}}}_{k,i,j}-{\mathbf{X}}_{k,i,j}\right)^{2}}.$
Another metric widely used in hyperspectral imaging is the Spectral Angle Mapper (SAM) (Van der Meer, 2006; Kuching, 2007; Park et al., 2007), although it has not yet found its way into the relevant computer vision literature. SAM emphasizes the spectral accuracy compared to the previous metrics, which are more forgiving of large errors in individual spectral channels:
(7) ${\mathrm{\mathbf{SAM}}}=\frac{1}{MN}\sum_{i,j}\cos^{-1}\left(\frac{\sum_{k}\hat{{\mathbf{X}}}_{k,i,j}{\mathbf{X}}_{k,i,j}}{\sqrt{\sum_{k}\hat{{\mathbf{X}}}_{k,i,j}^{2}\sum_{k}{\mathbf{X}}_{k,i,j}^{2}}}\right).$
@@
(8) ${\mathrm{\mathbf{PSNR}}}=\frac{1}{K}\sum_{k}20\log_{10}\left(\frac{\textrm{MAX}}{\sqrt{\textrm{MSE}_{k}}}\right),$
@@ 3.5. Training details
In our study, we conduct all the experiments across various datasets and network architectures. Following the methodologies proposed by the latest champion network MST++ (Cai et al., 2022b), we employ their patch-wise training approach (patches of 128$\times$128 pixels).
3.5.1. Data Preparation
@@
Throughout the experiments in this work, we follow the same data format (Matlab-compatible mat files) for hyperspectral images in the ARAD1K dataset (Arad et al., 2022). To be consistent, we also convert the raw hyperspectral datacubes in the CAVE (Yasuma et al., 2010), ICVL (Arad and Ben-Shahar, 2016), and KAUST (Li et al., 2021) datasets to this format. The data values are normalized by their respective bit-depths such that the data range is [0.0, 1.0]. The training and validation sets in ARAD1K are kept the same as offered in the NTIRE 2022 spectral recovery challenge (Arad et al., 2022), i.e., 900 files for training, and 50 for validation. We split the CAVE, ICVL, and KAUST datasets by 90% for training, and 10% for validation. Following the training strategy in MST++ (Cai et al., 2022b), we keep the training and validation lists fixed.
3.5.2. Metamer Generation
@@
We adopt the metameric black method (van Trigt, 1994; Finlayson and Morovic, 2005; Alsam and Lenz, 2007) to generate metamers from the original hyperspectral data. By varying the coefficient of the metameric black term, we could generate metamers that project to the same RGB color. Note that the original datacube corresponds to $\alpha=1$, and the fundamental metamer corresponds to $\alpha=0$. For the experiments with fixed metamers, we use the fundamental metamers to complement the original standard data. This is sufficient to demonstrate our findings. Other arbitrary values would result in the same conclusions. A more aggressive setting is to vary $\alpha$ as a variable to account for the infinite possible metamers in a more realistic situation.
3.5.3. Training and Validation Procedure
@@
Following the training strategies in MST++ (Cai et al., 2022b), we sub-sample the hyperspectral datacubes and the corresponding RGB images into overlapping patches of 128$\times$128. Spatial augmentations, such as random rotation, vertical flipping, and horizontal flipping are randomly applied to the training patches. In the validation step, we calculate the evaluation metrics (MRAE, RMSE, PSNR, and SAM) on the full spatial resolution for the ARAD1K (482$\times$512), CAVE (512$\times$512), and KAUST (512$\times$512) datasets. Note that this is different from MST++ (Cai et al., 2022b) where only the central 256$\times$256 regions are evaluated. The ICVL dataset has a very large spatial resolution (1300$\times$1392). To be consistent with other datasets, we evaluate ICVL only in the central 512$\times$512 regions.
@@
Similar as MST++, in each epoch, we train the networks for 1000 iterations, with a total number of 300 epochs. All the reported results are evaluated at the end of the training epochs, i.e., 300k iterations. We find that the training iterations are sufficient to achieve convergence in all our proposed experiments. Hyperparameters, such as learning rate and batch size, are tuned to achieve the best performance for each network on each dataset.
All the experiments are conducted on an NVIDIA A100 GPU (80 GB memory).
4. Finding 1: Atypical Overfitting
@@
Although it is well-known that deep neural networks may suffer from overfitting problems, we find that the overfitting behavior in spectral reconstruction is atypical and difficult to notice with standard evaluations. Here we introduce minimalist changes to the ARAD1K dataset used in NTIRE 2022 challenge (Arad et al., 2022) in three experiments to demonstrate it. We evaluate 11 top-performant neural network architectures, namely MST++ (Cai et al., 2022b), MST-L (Cai et al., 2022a), MPRNet (Zamir et al., 2021), Restormer (Zamir et al., 2022), MIRNet (Zamir et al., 2020), HINet (Chen et al., 2021), HDNet (Hu et al., 2022), AWAN (Li et al., 2020), EDSR (Lim et al., 2017), HRNet (Zhao et al., 2020), and HSCNN+ (Shi et al., 2018).
4.1. Training with less data
@@
First, we make a simple change to the training of the participating networks in the NTIRE 2022 challenge (Arad et al., 2022). While keeping all the training settings intact, we randomly choose only 50% or 20% of the original training data, respectively, to train the candidate networks and validate the performance on the original validation data. We illustrate the validation curves for MST++ (Cai et al., 2022b) in Fig. 3. See the Supplemental Material for the results of other networks. We summarize the results for 100% and 50% training data in Table 2 for all networks.
@@
Figure 3. Validation performance for MST++ (Cai et al., 2022b) with 100%, 50%, and 20% of the original training data on ARAD1K (Arad et al., 2022).
Although the performance with less training data deviates mildly in MRAE, RMSE, and PSNR, the spectral accuracy SAM (highlighted in bold in Table 2) is surprisingly less affected. In particular, some networks (e.g., MST++ (Cai et al., 2022b), MIRNet (Zamir et al., 2020)) achieve exactly the same SAM scores. MST-L (Cai et al., 2022a) (50%) even improves SAM slightly, placing itself the best among all.
Table 2. Performance comparison with 100% and 50% of training data for different methods on the original ARAD1K dataset.
@@
4.2. Validation with unseen data
Table 3. Evaluation of the pre-trained MST++ model on synthesized validation data from the ARAD1K dataset with different noise levels, compression quality, and realistic optical aberrations.
@@ 4.2. Validation with unseen data
(Arad et al., 2022). Only the spectrum-to-color projection was considered, and no aberrations of the optical system were simulated.
In our experiments, we generate new RGB images using the same methodology and calibration data, but different noise and compression settings. Specifically, we use the SRF data for a Basler ace 2 camera (model A2a5320-23ucBAS) known to the networks, and simulate Poisson noise at varying noise levels by controlling the number of photon electrons (npe). We adopt the same rudimentary in-camera image signal processing pipeline. As an illustrative example, we use MST++ (Cai et al., 2022b) in Table LABEL:tab:unseen, Row 1 as a reference for comparison; results for other networks can be found in the Supplemental Material.
@@
Second, in Row 3, we generate noiseless RGB images, but in lossless PNG format, as opposed to the JPEG (Q = 65) in Row 2. Note that JPEG compression is not necessary for the core inverse problem in hyperspectral imaging, since raw data could be readily obtained from the sensors. This results in paradoxical reconstruction performance. MRAE and SAM improve compared to Row 2 (but are still worse than Row 1), while RMSE and PSNR deteriorate further. Considering this only eliminates image compression, and the networks were trained on MRAE (Cai et al., 2022b), we can confirm that the network indeed overfits the specific unknown JPEG compression used in the challenge (Arad et al., 2022).
Third, we consider a more realistic imaging scenario, in which we eliminate the impact of unnecessary compression by employing the lossless PNG format to save the RGB images (equivalent to using raw camera data). We adopt moderate noise levels (npe = 1000) and realistic optical aberrations from a recent double Gauss lens patent (Ichimura, 2021) to mimic a real photographic camera. We can observe a further performance drop in Row 4, which provides additional evidence that the network overfits the unknown parameters in the image simulation pipeline (Arad et al., 2022). When used under realistic imaging conditions, the performance degrades significantly.
4.3. Cross-dataset validation
In addition, we inspect the effects of different datasets (cf. Table 1) on the performance. We train the MST++ network on the four datasets with the same image simulation parameters. To eliminate the impact of other factors, we choose the ideal noiseless and aberration-free condition without compression. In the validation, we use our trained model on ARAD1K dataset to validate on the other three datasets, respectively. In Table 4, we compare the performance with the models both trained and validated on the original datasets. Results for other networks can be found in the Supplement. They all illustrate the same difficulties in generalization.
Table 4. Cross-dataset validation example using MST++ (Cai et al., 2022b).
@@ 4.3. Cross-dataset validation
We also point out that the CAVE dataset (Yasuma et al., 2010), although smaller and older than the others, is more difficult to train for better performance. This is probably due to the fact that CAVE consists of several challenging scenes of real and fake objects that appear in similar colors, but other datasets comprise less aggressive natural scenes.
4.4. Discussion
The experiments conducted in this section clearly highlight several shortcomings with respect to the existing datasets: (1) they lack diversity in nuisance parameters such as noise and compression ratios, and (2) they lack scene diversity. While we show in this document the results for the largest available dataset (ARAD1K), the Supplement shows consistent results for all the datasets. Both of these aspects result in over-fitting and prevent the networks from learning the general spectral image restoration task. Next we specifically analyze the effect of metamerism; the analysis of the impact of optical aberrations will be deepened in Section 6.
@@ 4.4. Discussion
Figure 4. Validation with metamers for MST++ (Cai et al., 2022b). An example Scene ARAD_1K_0944 is shown to visualize the standard and metamer datacubes. Top left: the standard and metamer data result in similar color images. Bottom left: ground-truth and reconstructed spectra from two labeled points. Right: ground-truth and reconstructed spectral images in 420 nm, 500 nm, 550 nm, 580 nm, and 660 nm.
5. Finding 2: Metameric Failure
@@
5.1. Validation with metamers
We generate metamer datacubes (metamer data) from the original ARAD1K dataset (standard data) using the metameric black method (Finlayson and Morovic, 2005). For this set of experiments, we fix the coefficient $\alpha=0$ in Eq. (3). We choose a realistic imaging condition as used in Table LABEL:tab:unseen, Row 4, and keep it the same for both cases. The validation results on the ARAD1K dataset for existing pre-trained networks are summarized in Table 5. We also visualize the reconstructed spectral images in five arbitrary bands (420 nm, 500 nm, 550 nm, 580 nm, and 660 nm) and spectra of two points in Fig. 4 for Scene ARAD_1K_0944 from the validation set.
Table 5. Validation performance for different pre-trained models on standard (std) data and metamer (met) adversary synthesized from the ARAD1K dataset (Arad et al., 2022).
@@ 5.1. Validation with metamers
From the numerical results in Table 5, it is apparent that all the existing methods experience catastrophic performance drop in terms of MRAE and SAM in the presence of metamers, which we call metameric failure. The MRAE (cf. Eq. (4)) may yield large values when big errors occur for dark ground-truth pixels (see the exemplary spectra in Fig. 4). The SAM values become large when the spectra are essentially dissimilar with each other. RMSE and PSNR do not capture the spectral differences as well, since they average out differences in the spatial and spectral dimensions.
The visual results in Fig. 4 show that the reconstruction results are very close to each other for both standard and metamer data, because the input RGB images are quite similar. However, distinct differences exist in the scene for certain spectral bands, e.g., the intensities of the yellow and green parts of the slide (blue box) in 500 nm band vary in the standard data, but remain identical in the metamer data. The reconstructions fail to reflect this important difference. All spectral images are displayed on the same global intensity scale, so the brightness differences (green box) in corresponding images reflect the reconstruction artifacts for the metamer data.
5.2. Training with metamers

### v3

@@ III-C Modeling metamerism
Interestingly, metamer generation from existing spectra has been studied in color science and spectral rendering to accurately model the scenes using various methods, e.g., metameric black [60, 26, 3] and spectral uplifting [35, 58, 12]. In this work, to support our analysis, we adopt the metameric black approach to generate metamers, whereas we note that other metamer generation methods can also be employed for the same purpose.
A spectrum ${\mathbf{S}}$ can be projected onto two orthogonal subspaces, one for the fundamental metamer ${\mathbf{S}}^{*}$, and the other for metameric black ${\mathbf{B}}$ [20, 61]. A new metameric spectrum ${\mathbf{S}}^{\prime}$ is then
${\mathbf{S}}^{\prime}={\mathbf{S}}^{*}+\alpha{\mathbf{B}},$ (3)
@@
III-D Performance evaluation metrics
Consider a hyperspectral image ${\mathbf{X}}_{k,i,j}$ and its estimate $\hat{{\mathbf{X}}}_{k,i,j}$, where $k$ is the spectral index, and $i,j$ are spatial indices. The reconstruction quality can be evaluated in various ways. The NTIRE 2022 spectral reconstruction challenge [7] adopts two numerical metrics, Mean Relative Absolute Error (MRAE),
${\mathrm{\mathbf{MRAE}}}=\frac{1}{KMN}\sum_{k,i,j}\left|\hat{{\mathbf{X}}}_{k,i,j}-{\mathbf{X}}_{k,i,j}\right|/{\mathbf{X}}_{k,i,j},\$ (4)
@@ III-D Performance evaluation metrics
${\mathrm{\mathbf{RMSE}}}=\sqrt{\frac{1}{KMN}\sum_{k,i,j}\left(\hat{{\mathbf{X}}}_{k,i,j}-{\mathbf{X}}_{k,i,j}\right)^{2}}.$ (5)
Another metric widely used in hyperspectral imaging is the Spectral Angle Mapper (SAM) [59, 40, 48], although it has not yet found its way into the relevant computer vision literature. SAM emphasizes the spectral accuracy compared to the previous metrics, which are more forgiving of large errors in individual spectral channels:
${\mathrm{\mathbf{SAM}}}=\frac{1}{MN}\sum_{i,j}\cos^{-1}\left(\frac{\sum_{k}\hat{{\mathbf{X}}}_{k,i,j}{\mathbf{X}}_{k,i,j}}{\sqrt{\sum_{k}\hat{{\mathbf{X}}}_{k,i,j}^{2}\sum_{k}{\mathbf{X}}_{k,i,j}^{2}}}\right).$ (6)
@@
${\mathrm{\mathbf{PSNR}}}=\frac{1}{K}\sum_{k}20\log_{10}\left(\frac{\textrm{MAX}}{\sqrt{\textrm{MSE}_{k}}}\right),$ (7)
@@ III-E Training details
In our study, we conduct all the experiments across various datasets and network architectures. Following the methodologies proposed by the latest champion network MST++ [14], we employ their patch-wise training approach (patches of 128$\times$128 pixels).
III-E1 Data Preparation
@@
Throughout the experiments in this work, we follow the same data format (Matlab-compatible mat files) for hyperspectral images in the ARAD1K dataset [7]. To be consistent, we also convert the raw hyperspectral datacubes in the CAVE [67], ICVL [4], and KAUST [42] datasets to this format. The data values are normalized by their respective bit-depths such that the data range is [0.0, 1.0]. The training and validation sets in ARAD1K are kept the same as offered in the NTIRE 2022 spectral recovery challenge [7], i.e., 900 files for training, and 50 for validation. We split the CAVE, ICVL, and KAUST datasets by 90% for training, and 10% for validation. Following the training strategy in MST++ [14], we keep the training and validation lists fixed.
III-E2 Metamer Generation
@@
We adopt the metameric black method [60, 26, 3] to generate metamers from the original hyperspectral data. By varying the coefficient of the metameric black term, we could generate metamers that project to the same RGB color. Note that the original datacube corresponds to $\alpha=1$, and the fundamental metamer corresponds to $\alpha=0$. For the experiments with fixed metamers, we use the fundamental metamers to complement the original standard data. This is sufficient to demonstrate our findings. Other arbitrary values would result in the same conclusions. A more aggressive setting is to vary $\alpha$ as a variable to account for the infinite possible metamers in a more realistic situation.
III-E3 Training and Validation Procedure
@@
Following the training strategies in MST++ [14], we sub-sample the hyperspectral datacubes and the corresponding RGB images into overlapping patches of 128$\times$128. Spatial augmentations, such as random rotation, vertical flipping, and horizontal flipping are randomly applied to the training patches. In the validation step, we calculate the evaluation metrics (MRAE, RMSE, PSNR, and SAM) on the full spatial resolution for the ARAD1K (482$\times$512), CAVE (512$\times$512), and KAUST (512$\times$512) datasets. Note that this is different from MST++ [14] where only the central 256$\times$256 regions are evaluated. The ICVL dataset has a very large spatial resolution (1300$\times$1392). To be consistent with other datasets, we evaluate ICVL only in the central 512$\times$512 regions.
@@
Similar as MST++, in each epoch, we train the networks for 1000 iterations, with a total number of 300 epochs. All the reported results are evaluated at the end of the training epochs, i.e., 300k iterations. We find that the training iterations are sufficient to achieve convergence in all our proposed experiments. Hyperparameters, such as learning rate and batch size, are tuned to achieve the best performance for each network on each dataset. All the experiments are conducted on an NVIDIA A100 GPU (80 GB memory).
IV Finding 1: Atypical Overfitting
@@
Although it is well-known that deep neural networks may suffer from overfitting problems, we find that the overfitting behavior in spectral reconstruction is atypical and difficult to notice with standard evaluations. Here we introduce minimalist changes to the ARAD1K dataset used in NTIRE 2022 challenge [7] in three experiments to demonstrate it. We evaluate 11 top-performant neural network architectures, namely MST++ [14], MST-L [13], MPRNet [70], Restormer [68], MIRNet [69], HINet [17], HDNet [31], AWAN [41], EDSR [43], HRNet [72], and HSCNN+ [54].
IV-A Training with less data
@@
First, we make a simple change to the training of the participating networks in the NTIRE 2022 challenge [7]. While keeping all the training settings intact, we randomly choose only 50% or 20% of the original training data, respectively, to train the candidate networks and validate the performance on the original validation data. We illustrate the validation curves for MST++ [14] in Fig. 3. See the Supplemental Material for the results of other networks. We summarize the results for 100% and 50% training data in Table II for all networks.
@@
Figure 3: Validation performance for MST++ [14] with 100%, 50%, and 20% of the original training data on ARAD1K [7].
Although the performance with less training data deviates mildly in MRAE, RMSE, and PSNR, the spectral accuracy SAM (highlighted in bold in Table II) is surprisingly less affected. In particular, some networks (e.g., MST++ [14], MIRNet [69]) achieve exactly the same SAM scores. MST-L [13] (50%) even improves SAM slightly, placing itself the best among all.
TABLE II: Performance comparison with 100% and 50% of training data for different methods on the original ARAD1K dataset.
@@
IV-B Validation with unseen data
TABLE III: Evaluation of the pre-trained MST++ model on synthesized validation data from the ARAD1K dataset with different noise levels, compression quality, and realistic optical aberrations.
@@ IV-B Validation with unseen data
[7]. Only the spectrum-to-color projection was considered, and no aberrations of the optical system were simulated.
In our experiments, we generate new RGB images using the same methodology and calibration data, but different noise and compression settings. Specifically, we use the SRF data for a Basler ace 2 camera (model A2a5320-23ucBAS) known to the networks, and simulate Poisson noise at varying noise levels by controlling the number of photon electrons (npe). We adopt the same rudimentary in-camera image signal processing pipeline. As an illustrative example, we use MST++ [14] in Table LABEL:tab:unseen, Row 1 as a reference for comparison; results for other networks can be found in the Supplemental Material.
@@
Second, in Row 3, we generate noiseless RGB images, but in lossless PNG format, as opposed to the JPEG (Q = 65) in Row 2. Note that JPEG compression is not necessary for the core inverse problem in hyperspectral imaging, since raw data could be readily obtained from the sensors. This results in paradoxical reconstruction performance. MRAE and SAM improve compared to Row 2 (but are still worse than Row 1), while RMSE and PSNR deteriorate further. Considering this only eliminates image compression, and the networks were trained on MRAE [14], we can confirm that the network indeed overfits the specific unknown JPEG compression used in the challenge [7].
Third, we consider a more realistic imaging scenario, in which we eliminate the impact of unnecessary compression by employing the lossless PNG format to save the RGB images (equivalent to using raw camera data). We adopt moderate noise levels (npe = 1000) and realistic optical aberrations from a recent double Gauss lens patent [33] to mimic a real photographic camera. We can observe a further performance drop in Row 4, which provides additional evidence that the network overfits the unknown parameters in the image simulation pipeline [7]. When used under realistic imaging conditions, the performance degrades significantly.
IV-C Cross-dataset validation
In addition, we inspect the effects of different datasets (cf. Table I) on the performance. We train the MST++ network on the four datasets with the same image simulation parameters. To eliminate the impact of other factors, we choose the ideal noiseless and aberration-free condition without compression. In the validation, we use our trained model on ARAD1K dataset to validate on the other three datasets, respectively. In Table IV, we compare the performance with the models both trained and validated on the original datasets. Results for other networks can be found in the Supplement. They all illustrate the same difficulties in generalization.
TABLE IV: Cross-dataset validation example using MST++ [14].
@@ IV-C Cross-dataset validation
We also point out that the CAVE dataset [67], although smaller and older than the others, is more difficult to train for better performance. This is probably due to the fact that CAVE consists of several challenging scenes of real and fake objects that appear in similar colors, but other datasets comprise less aggressive natural scenes.
IV-D Discussion
The experiments conducted in this section clearly highlight several shortcomings with respect to the existing datasets: (1) they lack diversity in nuisance parameters such as noise and compression ratios, and (2) they lack scene diversity. While we show in this document the results for the largest available dataset (ARAD1K), the Supplement shows consistent results for all the datasets. Both of these aspects result in over-fitting and prevent the networks from learning the general spectral image restoration task. Next we specifically analyze the effect of metamerism; the analysis of the impact of optical aberrations will be deepened in Section VI.
@@ IV-D Discussion
Figure 4: Validation with metamers for MST++ [14]. An example Scene ARAD_1K_0944 is shown to visualize the standard and metamer datacubes. Top left: the standard and metamer data result in similar color images. Bottom left: ground-truth and reconstructed spectra from two labeled points. Right: ground-truth and reconstructed spectral images in 420 nm, 500 nm, 550 nm, 580 nm, and 660 nm.
V Finding 2: Metameric Failure
@@
V-A Validation with metamers
We generate metamer datacubes (metamer data) from the original ARAD1K dataset (standard data) using the metameric black method [26]. For this set of experiments, we fix the coefficient $\alpha=0$ in Eq. (3). We choose a realistic imaging condition as used in Table LABEL:tab:unseen, Row 4, and keep it the same for both cases. The validation results on the ARAD1K dataset for existing pre-trained networks are summarized in Table V. We also visualize the reconstructed spectral images in five arbitrary bands (420 nm, 500 nm, 550 nm, 580 nm, and 660 nm) and spectra of two points in Fig. 4 for Scene ARAD_1K_0944 from the validation set.
TABLE V: Validation performance for different pre-trained models on standard (std) data and metamer (met) adversary synthesized from the ARAD1K dataset [7].
@@ V-A Validation with metamers
From the numerical results in Table V, it is apparent that all the existing methods experience catastrophic performance drop in terms of MRAE and SAM in the presence of metamers, which we call metameric failure. The MRAE (cf. Eq. (4)) may yield large values when big errors occur for dark ground-truth pixels (see the exemplary spectra in Fig. 4). The SAM values become large when the spectra are essentially dissimilar with each other. RMSE and PSNR do not capture the spectral differences as well, since they average out differences in the spatial and spectral dimensions.
The visual results in Fig. 4 show that the reconstruction results are very close to each other for both standard and metamer data, because the input RGB images are quite similar. However, distinct differences exist in the scene for certain spectral bands, e.g., the intensities of the yellow and green parts of the slide (blue box) in 500 nm band vary in the standard data, but remain identical in the metamer data. The reconstructions fail to reflect this important difference. All spectral images are displayed on the same global intensity scale, so the brightness differences (green box) in corresponding images reflect the reconstruction artifacts.
V-B Training with metamers